# Neural-Driven Image Editing

### v1

@@ 1 Introduction
    Are neural signals reliable indicators of user intention in image editing?
@@
    If yes, what aspects of condition information or cognitive state are captured by multimodal neural signals?
@@
    Do neural signals and speech provide comparable guidance for image editing, and can they offer complementary information?
To answer these questions, we first construct L-Mind, a comprehensive multimodal dataset comprising 23,928 image pairs with synchronously collected EEG, functional near infrared spectroscopy (fNIRS) [33], photoplethysmography (PPG) [34], head motion, and speech signals from 12 participants conceiving image editing tasks. Captured using a wireless, lightweight BCI system that supports unconstrained head movements and speech [35], L-Mind offers higher ecological validity under natural real-world conditions and supports robust training of brain-supervised generative models.
Building on L-Mind, we also propose LoongX, a hands-free image editing approach that innovatively integrates the proposed multimodal neural signal fusion strategy with a diffusion transformer (DiT) to translate neural intent into image edits. Unlike prior single-modal methods, LoongX integrates EEG, fNIRS, PPG, and head motion signals, extracting explicit user intentions from EEG signals across multiple scalp regions, incorporating cognitive load and emotional valence data from fNIRS, and capturing stress and engagement indicators through PPG and motion signals. We introduce two new modules to manage diverse multimodal input: the cross-scale state space (CS3) encoder for robust feature extraction and the dynamic gated fusion (DGF) module for comprehensive multimodal integration. We further pretrain these encoders via contrastive learning using combined large-scale datasets [36] and L-Mind to align neural features with semantic text embeddings.
@@
Our main contributions are summarized as:
@@ 3.2 Data Preprocessing
Hemodynamic signals ($\Delta$HbO, $\Delta$HbR, and $\Delta$HbT, where $\Delta$HbT = $\Delta$HbO + $\Delta$HbR) are band-pass filtered (0.01–0.5 Hz) to isolate relevant neural responses, averaged per hemisphere to reflect task-related brain activity.
@@ 4.1 Cross-Scale State Space Encoding
A single modality input signal $\mathbf{S}\in\mathbb{R}^{C\times L_{0}}$ is fed into an N-layer (e.g., $N=5$ for EEG) adaptive average pooling (AAP) module (we set $d=64$ for EEG):
@@
Extracted embedding is computed as the concatenation of the feature pyramid $\mathbf{P}=\operatorname{Concat}(\{\mathbf{P}_{i}\})$.
@@
[…]
@@
where $\mathbf{e}(t)$ denotes the latent state at time $t$, and $\hat{\mathbf{A}}$, $\hat{\mathbf{B}}$, $\hat{\mathbf{C}}$, $\hat{\mathbf{D}}$ are diagonal matrices that parameterize state transitions, input injection, state-to-output mapping, and direct input-to-output mapping, respectively. Due to the diagonal parameterization, the S3M block admits efficient computation with linear complexity $\mathcal{O}(L\log L)$. Through the S3M blocks, $\mathbf{Z}^{(1)}$ is down-sampled from length $L$ to $d_{m}$, yielding $\tilde{\mathbf{Z}}^{(1)}\!\in\!\mathbb{R}^{C\times d_{m}}$; $\mathbf{Z}^{(2)}$ is further permuted and down-sampled to length $d_{p}$ via an AAP, giving $\tilde{\mathbf{Z}}^{(2)}\!\in\!\mathbb{R}^{d_{m^{\prime}}\times d_{p}}$, where $d_{m^{\prime}}$ is set equal to $C$.
@@
$\mathbf{F}=\mathrm{ANP}\big{(}\operatorname{cat}_{c}(\tilde{\mathbf{Z}}^{(1)},\mathbf{P},\tilde{\mathbf{Z}}^{(2)})\in\mathbb{R}^{C\times L^{\prime}}\big{)}.$ (5)
where $L^{\prime}=d_{m}+d_{p}+\Sigma^{N}_{i}2^{i}\cdot d$. The concatenated feature was projected via Adaptive Nonlinear Projection (ANP), which consists of two fully-connected layers, LayerNorm, ReLU, and dropout. The final embeded feature $\mathbf{F}\in\mathbb{R}^{C^{\prime}\times d^{\prime}}$ is obtained.
@@ 4.2 Dynamic Gated Multimodal Fusion
We calculate instance-wise and layer-wise mean $\mu$ and variance $\sigma$ from input content embedding (e.g., EEG) $\mathbf{X}\!\in\!\mathbb{R}^{C\times L}$ and condition embedding (e.g., PPG) $\mathbf{Y}\!\in\!\mathbb{R}^{C\times L}$ for further fusion into $\tilde{\mathbf{H}}\!\in\!\mathbb{R}^{C\times L}$ while emphasising informative channels and suppressing noise:
$\underbrace{\begin{bmatrix}\mu_{\text{inst}}&\sigma_{\text{inst}}\\[4.0pt] \mu_{\text{layer}}&\sigma_{\text{layer}}\end{bmatrix}}_{\text{each entry }\in\mathbb{R}^{C\times 1}}=\begin{bmatrix}\dfrac{1}{L}\displaystyle\sum_{t}\mathbf{X}_{:,t}&\sqrt{\dfrac{1}{L}\displaystyle\sum_{t}\bigl{(}\mathbf{X}_{:,t}-\mu_{\text{inst}}\bigr{)}^{2}+\varepsilon}\\[10.0pt] \dfrac{1}{CL}\displaystyle\sum_{c,t}\mathbf{X}_{c,t}\,\mathbf{1}_{C}&\sqrt{\dfrac{1}{CL}\displaystyle\sum_{c,t}\bigl{(}\mathbf{X}_{c,t}-\mu_{\text{layer}}\bigr{)}^{2}+\varepsilon}\end{bmatrix},\qquad\varepsilon=10^{-3}.$ (6)
@@ 4.3 Conditional Diffusion
Specifically, DiT predicts a velocity $\boldsymbol{v}_{\theta}(\mathbf{I}_{t},\,t,\,\tilde{\mathbf{h}})$ that is used to iteratively refine the latent image in $T$ uniform steps,
@@ 5 Experiment
To answer each research question (RQ) raised in Sec. 1, we comprehensively evaluate LoongX’s capability of performing neural-driven image editing on the test set of L-Mind. The experimental setup, metrics from [51], and more implementation details are described in Sec. A.4.1. We choose OminiControl [49] as our baseline as it supports the text-conditioned image-editing based on DiTs.
@@ 5.1 Reliability of Neural Signals
Answering RQ1: As shown in Table 1, neural signals can work as reliable indicators to drive semantic and structural image editing, outperforming language-only baselines on key perceptual metrics. Remarkably, neural-signal-only LoongX surpasses the text-based OminiControl baseline in discriminability (CLIP-I: $0.6605$ vs. $0.6558$) and robustness (DINO: $0.4812$ vs. $0.4636$). This demonstrates neural signals’ intrinsic capability to carry reliable semantic information for image editing. While pixel-based metrics (L1/L2) show slight increases, this trade-off highlights the prioritization of semantic fidelity over pixel-level stickiness. Introducing speech cues fused with neural signals further enhances semantic alignment (CLIP-T reaching the highest $0.2588$), underscoring their combined strength in capturing nuanced user intentions for hands-free image editing.
@@ 5.2 Ablation Studies on Modality Contribution
Answering RQ2: Different neural signal modalities contribute complementary strengths, enhancing discriminability, robustness, and semantic precision, respectively. We examine modality contributions in detail, as shown in Fig. 5. EEG signals alone enable basic high-level semantic editing, supported by the semantic discriminability of the extracted features (CLIP-I: $0.5457$). Integrating fNIRS significantly improves feature robustness (DINO: from $0.2963$ to $0.4811$), highlighting the complementary nature of hemodynamic responses in enhancing signal completeness and structural fidelity. Including PPG and Motion improves global physiological awareness and indicates sensitivity to subtle engagement patterns (e.g., heartbeat and user movements) that express editing intent. They both contribute to features’ robustness and completeness to ensure stable CLIP-T score gains.
Fig. 5 further explores EEG channels’ individual contributions, where each channel corresponds to a specific scalp region as detailed in Table 3. The occipital cortex channel (Oz), which is in charge of visual processing, emerges as dominant in global editing effect (CLIP-I: $0.6619$) and robustness (DINO: $0.4873$), affirming its critical role in basic visual perception and processing tasks. Conversely, the frontopolar cortex (Fpz) provides superior semantic alignment (CLIP-T: $0.2481$), consistent with its association with more complex cognitive processes. Specifically, Fpz provides decision control and attention regulation compared with basic visual perception provided by Oz, which precisely confirms the discovery patterns in medical anatomy. This channel-specific analysis provides insights valuable for targeted applications or constrained hardware settings.
Figure 4: Evaluation results for different signal combinations using DGF.
@@
Figure 6: Breakdown results of text and neural-driven image editing metrics. BG: background.
@@
Answering RQ3: Neural signals excel in low-level visual edits, while language excels in high-level semantics; combining both yields the most effective hybrid control. As shown in Fig. 6, neural signals (N) are particularly effective for more intuitive tasks like global texture editing (higher CLIP-I), reflecting strong visual discriminability and structural coherence. In object editing, neural signals are more capable of object removal than others, demonstrating their strength in conveying intuitive intent, though they remain limited in handling complex semantics. In contrast, text instructions (T) are inherently stronger in high-level semantic tasks (e.g., restore), highlighting their advantage in semantic alignment. When combined, neural signals and speech (N+S) inputs achieve the best semantic alignment (CLIP-T: $0.2588$), showcasing the superior effectiveness of hybrid conditioning in capturing complex user intentions.
@@ 5.4 Ablation Studies on Model Architecture
Each architectural component of LoongX contributes uniquely, and their composition, especially with pretraining, unlocks comprehensive performance potential. The ablation study in Table 2 is conducted under the setting of fusing all signals and speech, exploring the impact of each proposed module. CS3 encoder basically enhances the completeness and smoothness of features via extracted features, reducing pixel-level error (L2 decreased by $5\%$), and DGF primarily enhances semantic alignment with textual instructions (CLIP-T improvement: $3.5\%$). Supplemented by pre-training, LoongX reaches optimal performance, indicating the important role of robust multimodal alignment and structured representation learning in maximizing editing performance.
@@ 5.5 Qualitative Analysis and Limitations
Qualitative examples confirm LoongX’s intuitive editing capabilities, with limitations mostly arising in abstract or ambiguous complex intentions. Qualitative results in Fig. 7 illustrate that neural-signal-driven editing effectively handles visual and structural modifications, such as background replacing and global adjustments. However, the fused neural-language method better captures nuanced instructions involving abstract semantics (e.g., “modify the text information"). Despite significant advancements, entity consistency (e.g., the style of the little girl in Fig. 7(b)) remains a limitation of current editing models. Moreover, highly abstract or ambiguous instructions occasionally still pose challenges (e.g., “winged white animal in Fig 11" and several more failure cases shown in Fig. 14), indicating areas where further refinement of entity interpretation and disambiguation from neural data is necessary.
@@ 6 Conclusion
We presented LoongX, a novel framework for hands-free image editing by conditioning diffusion models on multimodal neural signals, achieving comparable or superior performance to traditional text-driven baselines. Looking forward, the portability of our wireless setup opens exciting possibilities for real-world applications in immersive environments. Future work may explore integrating LoongX with VR/XR platforms to enable intuitive cognitive interaction, and further align neural representations with world models for projecting human intention into an interactive virtual world, paving the way toward mind-driven control in fully synthetic realities.
@@ A.1 Preliminary: Analysis of Brain Region Function and Mechanism
Table 3: Functional Roles of Multimodal Neural Signals in Hands-Free Image Editing. The Ch refers to the channel.
@@
To enable hands-free image editing, we integrate multimodal neural and physiological signals to decode user intent and cognitive state in real time. Specifically: Midline parietal EEG (Pz) reflects spatial attention and visuomotor integration, supporting the allocation of attention to target areas and coordination of motor plans during editing tasks [58]. Right prefrontal EEG (Fp2) is linked to emotional regulation and motivational drive, aligning with findings from frontal alpha asymmetry studies [59]. Frontopolar EEG (Fz/Fpz) tracks attention control and task initiation, facilitating the onset and modulation of editing operations [60, 61]. Occipital EEG (Oz) encodes visual perception and image processing load, enabling evaluation of visual changes and edit quality [62]. fNIRS over anterior medial prefrontal cortex (aMPFC) measures cognitive load, motivation, and emotional valence through hemodynamic activity [63]. PPG signals from the same region capture heart rate variability and autonomic arousal, which reflect user stress and engagement [31, 32]. This multimodal mapping (Fig. 8 (b)) enables the system to continuously adapt to users’ cognitive focus, affective state, and mental workload, thereby supporting a more responsive and intuitive editing experience.

### v2

@@ 1 Introduction
    Can neural signals alone drive instruction-based image editing?
@@
    If yes, what kind of information do multimodal neural signals contribute?
@@
    How do neural-signal conditions compare and complement natural-language instructions?
To answer these questions, we construct L-Mind, a comprehensive multimodal dataset comprising 23,928 image pairs with synchronously collected EEG, functional near infrared spectroscopy (fNIRS) [33], photoplethysmography (PPG) [34], head motion, and speech signals from 12 participants conceiving image editing tasks. Captured using a wireless, lightweight BCI system that supports unconstrained head movements and speech [35], L-Mind offers higher ecological validity under natural real-world conditions and supports robust training of brain-supervised generative models.
Building on L-Mind, we also propose LoongX, a hands-free image editing approach that innovatively integrates the proposed multimodal neural signal fusion strategy with a diffusion transformer (DiT) to translate neural intent into image edits. Unlike prior single-modal methods, LoongX integrates EEG, fNIRS, PPG, and head motion signals, extracting explicit user intentions from EEG signals across multiple scalp regions, incorporating cognitive load and emotional valence data from fNIRS, and capturing stress and engagement indicators through PPG and motion signals. We introduce two new modules to manage diverse multimodal input: a cross-scale state space (CS3) encoder for robust feature extraction and a dynamic gated fusion (DGF) module for comprehensive multimodal integration. These encoders are pretrained via contrastive learning on combined large-scale datasets [36] and L-Mind to align neural features with semantic text embeddings.
@@
Our main contributions are summarized as follows:
@@ 3.2 Data Preprocessing
[…]
@@ 4.1 Cross-Scale State Space Encoding
A single modality input signal $\mathbf{S}\in\mathbb{R}^{C\times L_{0}}$ is fed into an $N$-layer adaptive average pooling (AAP) module:
@@
where we set $N=5$ and $d=64$ for EEG. The extracted embedding is computed as the concatenation of the feature pyramid $\mathbf{P}=\operatorname{Concat}(\{\mathbf{P}_{i}\})$.
@@
Specifically, the input signal $\mathbf{S}$ is padded from length $L_{0}$ to $L$, where $\mathbf{S}_{\mathrm{pad}}\in\mathbb{R}^{C\times L}$ with signal intensity normalized to [1,1][-1,1]. The padded signals and its permuted version $\mathbf{S}_{\mathrm{pm}}\in\mathbb{R}^{L\times C}$ are passed to two parallel S3M blocks, $\mathrm{S3M}^{(1)}$ and $\mathrm{S3M}^{(2)}$, respectively:
@@
where $\mathbf{e}(t)$ denotes the latent state at time $t$, and $\hat{\mathbf{A}}$, $\hat{\mathbf{B}}$, $\hat{\mathbf{C}}$, $\hat{\mathbf{D}}$ are diagonal matrices that parameterize state transitions, input injection, state-to-output mapping, and direct input-to-output mapping, respectively. Due to the diagonal parameterization, the S3M block admits efficient computation with linear complexity $\mathcal{O}(L\log L)$. Through the S3M blocks, $\mathbf{Z}^{(1)}$ is down-sampled from length $L$ to $d_{m}$, yielding $\tilde{\mathbf{Z}}^{(1)}\!\in\!\mathbb{R}^{C\times d_{m}}$; $\mathbf{Z}^{(2)}$ is permuted and down-sampled via an AAP, giving $\tilde{\mathbf{Z}}^{(2)}\!\in\!\mathbb{R}^{C\times d_{p}}$.
@@
$\mathbf{F}=\mathrm{ANP}\big{(}\operatorname{cat}_{c}(\tilde{\mathbf{Z}}^{(1)},\mathbf{P},\tilde{\mathbf{Z}}^{(2)})\in\mathbb{R}^{C\times d^{\prime}}\big{)}.$ (5)
where $d^{\prime}=d_{m}+d_{p}+\Sigma^{N}_{i}2^{i}\cdot d$. The concatenated feature was projected via Adaptive Nonlinear Projection (ANP), which consists of two fully-connected layers, LayerNorm, ReLU, and dropout. The final embeded feature $\mathbf{F}\in\mathbb{R}^{C^{\prime}\times L^{\prime}}$ is obtained.
@@ 4.2 Dynamic Gated Multimodal Fusion
We calculate instance-wise and layer-wise mean $\mu$ and variance $\sigma$ from input content embedding $\mathbf{X}\!\in\!\mathbb{R}^{C^{\prime}\times L^{\prime}_{X}}$ and condition embedding $\mathbf{Y}\!\in\!\mathbb{R}^{C^{\prime}\times L^{\prime}_{Y}}$ for further fusion into $\tilde{\mathbf{H}}\!\in\!\mathbb{R}^{C^{\prime}\times L^{\prime}_{X}}$ while emphasising informative channels and suppressing noise:
$\underbrace{\begin{bmatrix}\mu_{\text{inst}}&\sigma_{\text{inst}}\\[4.0pt] \mu_{\text{layer}}&\sigma_{\text{layer}}\end{bmatrix}}_{\text{each entry }\in\mathbb{R}^{C\times 1}}=\begin{bmatrix}\dfrac{1}{L_{X}^{{}^{\prime}}}\displaystyle\sum_{t}\mathbf{X}_{:,t}&\sqrt{\dfrac{1}{L_{X}^{{}^{\prime}}}\displaystyle\sum_{t}\bigl{(}\mathbf{X}_{:,t}-\mu_{\text{inst}}\bigr{)}^{2}+\varepsilon}\\[10.0pt] \dfrac{1}{C^{\prime}L_{X}^{{}^{\prime}}}\displaystyle\sum_{c,t}\mathbf{X}_{c,t}\,\mathbf{1}_{C^{\prime}}&\sqrt{\dfrac{1}{C^{\prime}L_{X}^{{}^{\prime}}}\displaystyle\sum_{c,t}\bigl{(}\mathbf{X}_{c,t}-\mu_{\text{layer}}\bigr{)}^{2}+\varepsilon}\end{bmatrix},\qquad\varepsilon=10^{-3}.$ (6)
@@ 4.3 Conditional Diffusion
[…]
@@ 5 Experiment
To answer the research questions (RQs) asked in Sec. 1, we conduct a comprehensive evaluation of LoongX’s neural-driven image editing capabilities on the L-Mind test set. The experimental setup, error metrics [51], and implementation details are described in Sec. A.4.1. We choose OminiControl [49] as our baseline, which supports the text-conditioned image-editing based on DiTs.
@@ 5.1 Reliability of Neural Signals
Answer RQ1: Neural signals can serve as reliable indicators to drive image editing, outperforming text-instructed baselines on key metrics.
As shown in Table 1, neural-signal-only LoongX outperforms the text-based OminiControl baseline in semantic discriminability (CLIP-I: $0.6605$ vs. $0.6558$) and robustness (DINO: $0.4812$ vs. $0.4636$), which demonstrates the potential of neural signals as a standalone modality that carries semantic information for image editing. The slightly higher L1 and L2 errors indicate better preservation of semantic fidelity over pixel-level accuracy. Combining speech cues with neural signals boosts semantic alignment, reaching a peak CLIP-T score of 0.2588 and demonstrating their joint effectiveness in capturing nuanced user intentions for hands-free image editing.
@@ 5.2 Ablation Studies on Modality Contribution
Answer RQ2: Different neural signal modalities contribute complementary strengths, enhancing discriminability, robustness, and semantic precision, respectively.
Modality contributions are compared in Fig. 5. EEG signals alone enable basic high-level semantic editing, supported by the semantic discriminability of the extracted features (CLIP-I: $0.5457$). Integrating fNIRS significantly improves feature robustness (DINO: from $0.2963$ to $0.4811$), highlighting the complementary nature of hemodynamic responses in enhancing signal completeness and structural fidelity. Including PPG and Motion improves global physiological awareness and indicates sensitivity to subtle engagement patterns (e.g., heart rate and user movements) that express editing intent. They both contribute to features’ robustness and completeness to ensure stable CLIP-T score gains.
We show contribution of each individual EEG channel in Fig. 5, where each channel corresponds to a specific scalp region as detailed in Table 3. The occipital cortex channel (Oz), which is in charge of visual processing, emerges as dominant in global editing effect (CLIP-I: $0.6619$) and robustness (DINO: $0.4873$) to finer details, affirming its critical role in basic visual perception and processing tasks. Conversely, the frontopolar cortex (Fpz) provides superior semantic alignment (CLIP-T: $0.2481$), consistent with its association with more complex cognitive processes. Specifically, Fpz provides decision control and attention regulation compared with basic visual perception provided by Oz, which precisely confirms the discovery patterns in medical anatomy. This channel-specific analysis provides insights valuable for targeted applications or constrained hardware settings.
Figure 4: Evaluation of different signals combined by our DGF module.
@@
Answer RQ3: Neural signals excel in low-level visual edits, while language excels in high-level semantics; combining both yields comprehensive and optimal control.
Figure 6: Breakdown results of text and neural-driven image editing. BG: background.
The analysis of text and neural-driven image editing are shown in Fig. 6, neural signals (N) demonstrate superior performance in global texture editing, with higher CLIP-I scores highlighting their strong visual and structural consistency. Neural signals also outperform other modalities in object editings, (e.g.,object removal), reflecting their strength in conveying intuitive intent, though they remain limited in handling complex semantics. Text instructions (T) are inherently stronger in high-level semantic tasks (e.g., image restoration), which indiciates their advantage in semantic alignment. Combined neural and speech (N+S) signals achieve the highest semantic alignment (CLIP-T: $0.2588$), showcasing the superior effectiveness of hybrid conditioning in capturing complex user intentions.
@@ 5.4 Ablation Studies on Model Architecture
Finding 1: Each component in the LoongX architecture contributes uniquely, and their joint pretraining through multimodal alignment maximizes their collective strengths, leading to improved editing accuracy across metrics.
The ablation study in Table 2 is conducted using the setting of fusing all signals and speech, exploring the impact of each proposed module. CS3 encoder enhances the completeness and smoothness of features via extracted features, reducing pixel-level error (L2 decreased by $5\%$), and DGF primarily improves semantic alignment with textual instructions (CLIP-T improvement: $3.5\%$). Supplemented by pre-training, LoongX reaches optimal performance, indicating the important role of robust multimodal alignment and sementic representation learning in maximizing editing performance.
@@ 5.5 Qualitative Analysis and Limitations
Finding 2: LoongX demonstrates image editing capabilities from neural signals, though performance may vary slightly with complex or ambiguous intentions.
Qualitative results shown in Fig. 7 illustrate that neural-signal-driven editing effectively handles visual and structural modifications, such as background replacing and global adjustments. However, the fused neural-language method better captures nuanced instructions involving abstract semantics (e.g., “modify the text information"). Despite significant advancements, entity consistency (e.g., the style of the little girl in Fig. 7(b)) remains a limitation of current editing models. Moreover, highly abstract or ambiguous instructions occasionally still pose challenges (e.g., “winged white animal in Fig 11" and several more failure cases shown in Fig. 14), indicating areas where further refinement of entity interpretation and disambiguation from neural data is necessary.
@@ 6 Conclusion
We presented LoongX, a novel framework for hands-free image editing by conditioning diffusion models on multimodal neural signals, achieving performance comparable to or superior to traditional text-driven baselines. Looking forward, the portability of our wireless setup opens exciting possibilities for real-world applications in immersive environments. Future directions include integrating LoongX with VR/XR to support intuitive cognitive interaction and aligning neural representations with world models for mind-driven control in virtual environments.
@@ A.1 Preliminary: Analysis of Brain Region Function and Mechanism
Table 3: Functional Roles of Multimodal Neural Signals in Hands-Free Image Editing.
@@
To enable hands-free image editing, we integrate multimodal neural and physiological signals to decode user intent and cognitive state in real time. The functional roles of multimodal neural signals are provided in Table.3. Specifically, midline parietal EEG (Pz) reflects spatial attention and visuomotor integration, supporting the allocation of attention to target areas and coordination of motor plans during editing tasks [58]. Right prefrontal EEG (Fp2) is linked to emotional regulation and motivational drive, aligning with findings from frontal alpha asymmetry studies [59]. Frontopolar EEG (Fz/Fpz) tracks attention control and task initiation, facilitating the onset and modulation of editing operations [60, 61]. Occipital EEG (Oz) encodes visual perception and image processing load, enabling evaluation of visual changes and edit quality [62]. fNIRS over anterior medial prefrontal cortex (aMPFC) measures cognitive load, motivation, and emotional valence through hemodynamic activity [63]. PPG signals from the same region capture heart rate variability and autonomic arousal, which reflect user stress and engagement [31, 32]. This multimodal mapping (Fig. 8 (b)) enables the system to continuously adapt to users’ cognitive focus, affective state, and mental workload, thereby supporting a more responsive and intuitive editing experience.